For the following equation, state the value of the discriminant and then describe the nature of the solutions. 17x^2-10x-20=0
Answer by stanbon(75887)

For the following equation, state the value of the discriminant and then describe the nature of the solutions. 17x^2-10x-20=0
------------------------------------
discriminant = b^2-4ac = 10^2 - 4*17*-20 = 100+1360 >0
---
Since the discriminat is > 0, there are two
unequal Real Number solutions.
===============================
Cheers,
Stan H.
===============================